OpenStudy (megannicole51):

I need an easier way to solve this problem than what my professor taught me please!! Exam tomorrow!!! How many terms of the Taylor series for ln(1+x) centered at x=0 do you need to estimate the value of ln(1.4) to three decimal points. I know ln(1+x)=x-x^2/2+x^3/3-... ln(1+.4) abs of En(.4) but thats all I know and understand

OpenStudy (megannicole51):

OH and f(.04)=ln(1.4)

OpenStudy (anonymous):

if i am not mistaken, find the term that give a number less than \(.001\) since the terms alternate, once you are there, the estimate is good to that many decimal places

OpenStudy (megannicole51):

wait what?

OpenStudy (anonymous):

DUde you can get a ti 89 titanium like mines and it will do it ;)

OpenStudy (anonymous):

the series alternates, and the terms get smaller and smaller, so the error can be no more than the last term

OpenStudy (megannicole51):

im studying for my exam and were not allowed to use any calculators during exams or quizzes so i dont use mine anymore to practice not using it

OpenStudy (anonymous):

\(\frac{.4^3}{3}=.021...\) so you need to go further out

OpenStudy (anonymous):

if it's centered at 0 is it now the maclurian series?

OpenStudy (megannicole51):

my professor has been using these weird formulas calculating the error and he's not explaining anything and the exam is tomorrow so I'm SOL

OpenStudy (anonymous):

hmm it is numerical exercise, can't think of a way to do it without checking the numbers. i mean if it was \(\ln(1.1)\) then the answer would be different if it was \(\ln(1)\) then the answer would be 1 term

OpenStudy (megannicole51):

what i need is steps to finding this problem just so i can get by with my exam please

OpenStudy (anonymous):

https://www.khanacademy.org/math/calculus/sequences_series_approx_calc/maclaurin_taylor/v/maclauren-and-taylor-series-intuition Watch the vids in this series. ANd just plug in values :D

OpenStudy (megannicole51):

this is using error not taylor polynomials

OpenStudy (anonymous):

try this